# The DMA Streaming Framework: Kernel-Level Buffer Orchestration for High-Performance AI Data Paths

### 26 Feb 2026 v1

@@ 1. Introduction
Modern AI systems are limited by host side data movement as often as by accelerator compute. Large language model workloads increasingly rely on point to point transfers between GPUs across hosts, including disaggregated inference where a prefill GPU produces KV cache state and a decode GPU consumes it. Transport systems can move bytes efficiently, but they commonly assume that buffers are already allocated on the correct NUMA node, shareable across devices without copies, registered for RDMA, and safe under completion pressure and teardown. These preconditions are not transport. They are buffer orchestration.
@@ 1.2. Motivating workloads beyond disaggregated inference
Disaggregated inference is a primary motivation because it requires point to point KV cache movement with receiver notification and bounded completion pressure. The same buffer orchestration layer is required in several adjacent workloads.
@@ 2.1. Kernel primitives used by dmaplane
Kernel verbs RDMA. The InfiniBand verbs API in ib_core provides a provider independent interface for RDMA. Kernel consumers allocate PDs, CQs, QPs, and register memory regions. Kernel space ownership is required when the memory being registered is not userspace memory.
@@ 2.2. Positioning and closest neighbors
NCCL. NCCL (NVIDIA Corporation, 2026a) implements GPU optimized collective communication and includes internal buffer pool and topology logic. It targets barrier synchronized collectives. dmaplane targets point to point pipelines and provides the kernel level buffer lifecycle and completion safety substrate beneath collective scheduling.
@@
nvidia-peermem. nvidia-peermem (NVIDIA Corporation, 2026b) bridges GPU BAR pages into the RDMA stack by registering them as memory regions. It addresses a narrow GPU to NIC sharing path. dmaplane incorporates an equivalent pinning and teardown discipline as one subsystem and extends it with NUMA placement, dma-buf export, flow control, and observability.
@@ 6.3. GPU memory access tiers
Derived: 64 KB / 3.2 ms = ~20 MB/s effective for the loopback entry, based on the
measured 64 KB loopback latency.
@@ 6.5.3. Not validated on hardware RDMA
The paper does not include performance measurements on hardware RDMA NICs such as ConnectX or EFA. GPU pinning uses the same peer to peer API used in GPUDirect RDMA environments, but NIC BAR DMA tiers are not measured in this work.
@@ C.1. Test suite map
Table C.1. Test suite map by subsystem.

### 25 Mar 2026 v2

@@ 1. Introduction
Modern AI systems are limited by host side data movement as often as by accelerator compute. Large language model workloads increasingly rely on point to point transfers between GPUs across hosts, including disaggregated inference where a prefill GPU produces KV cache state and a decode GPU consumes it (Patel et al., 2024; Zhong et al., 2024). Transport systems can move bytes efficiently, but they commonly assume that buffers are already allocated on the correct NUMA node, shareable across devices without copies, registered for RDMA, and safe under completion pressure and teardown. These preconditions are not transport. They are buffer orchestration.
@@ 1.2. Motivating workloads beyond disaggregated inference
Disaggregated inference (Patel et al., 2024; Zhong et al., 2024; Qin et al., 2024) is a primary motivation because it requires point to point KV cache movement with receiver notification and bounded completion pressure. The same buffer orchestration layer is required in several adjacent workloads.
@@ 2.1. Kernel primitives used by dmaplane
Kernel verbs RDMA. The InfiniBand verbs API in ib_core provides a provider independent interface for RDMA (Corbet et al., 2005). Kernel consumers allocate PDs, CQs, QPs, and register memory regions. Kernel space ownership is required when the memory being registered is not userspace memory.
@@ 2.2. Positioning and closest neighbors
NCCL. NCCL (NVIDIA Corporation, 2026b) implements GPU optimized collective communication and includes internal buffer pool and topology logic. It targets barrier synchronized collectives. dmaplane targets point to point pipelines and provides the kernel level buffer lifecycle and completion safety substrate beneath collective scheduling.
@@
nvidia-peermem. nvidia-peermem (NVIDIA Corporation, 2026c) bridges GPU BAR pages into the RDMA stack by registering them as memory regions. It addresses a narrow GPU to NIC sharing path. dmaplane incorporates an equivalent pinning and teardown discipline as one subsystem and extends it with NUMA placement, dma-buf export, flow control, and observability.
@@ 6.3. GPU memory access tiers
Note on GPU RDMA loopback entry: The ~20 MB/s effective rate is derived from the measured 64 KB loopback latency of 3.2 ms (64 KB / 3.2 ms $\approx$ 20 MB/s).
@@ 6.5.3. Not validated on hardware RDMA
The paper does not include performance measurements on hardware RDMA NICs such as ConnectX or EFA. GPU pinning uses the same peer to peer API used in GPUDirect RDMA environments (NVIDIA Corporation, 2026a), but NIC BAR DMA tiers are not measured in this work.
@@ C.1. Test suite map
Table 6. Test suite map by subsystem.